# Dynamic similarity of vortex shedding in a superfluid flowing past a penetrable obstacle

### v1

@@ IV CONCLUSION AND OUTLOOK
We have shown that vortex shedding past a penetrable obstacle in a two-dimensional superfluid can be organized by a flow-defined length scale, the effective diameter $D_{\mathrm{eff}}$. Defined as the transverse extent of the region where the compressible component of the superflow exceeds the local speed of sound, $D_{\mathrm{eff}}$ delineates the Mach-1 contour that marks the onset of compressibility-driven instability. $D_{\mathrm{eff}}$ expands with $v_{0}$ and quantitatively matches the observed width of the vorticity distribution, confirming that the dynamically relevant scale is that felt by the flow rather than the geometric obstacle size. When forces and frequencies are normalized by $D_{\mathrm{eff}}$, the Strouhal and drag coefficients collapse onto universal curves versus the superfluid Reynolds number $\mathrm{Re}_{s}$, with a clear transition at $\mathrm{Re}_{s}\approx 2$. Thus, $D_{\mathrm{eff}}$ provides a unified similarity length governing both the shedding dynamics and mean dissipation in penetrable obstacle flows. This finding provides a basis for extending the universality observed in classical fluid dynamics to regimes beyond any classical counterpart.
We have also examined the cases of impenetrable obstacles and non-Gaussian penetrable obstacles, showing that qualitative behaviors remain consistent with the present universal trend (see Appendix C). These results indicate that the effective-diameter framework can extend to a broader class of flow geometries beyond simple Gaussian profiles. However, it should be noted that the present framework, based on the existence of a supersonic interface defined by the Mach-1 contour, cannot be applied to cases with higher drive velocities $v_{0}/c_{\mathrm{s0}}>1$, as the entire flow field becomes supersonic and no bounded nucleation zone can be identified. In the fully supersonic regime, dissipation is expected to be governed by different dynamical mechanisms such as shock waves rather than vortex shedding [30, 55, 57], implying the need for a different characteristic length scale. In classical compressible flows, the shock layer sets the pressure drag for flow past a bluff body [56], indicating that a scale based on the shock layer is a natural choice in this limit.
@@
Finally, we note that the universality revealed here can be verified experimentally. The shedding frequency or $\mathrm{St}$ can be measured directly using a conventional time-of-flight image, as shown in Refs. [19, 27, 29]. In addition, the drag force can be measured without direct mechanical probes using a control-volume analysis, as suggested in Ref. [26]. Moreover, a recent proposal based on matter wave optics [58] suggests that both the superfluid velocity fields and the phase of the order parameter can be measured using phase–imprinting techniques [59, 60, 61], making it feasible to visualize the Mach-1 contour and measure $D_{\mathrm{eff}}$ directly. Such experimental methods would allow testing the predicted universal behaviors.
V acknowlegements
@@
We thank Haneul Kwak for providing the simulation code, Myeonghyeon Kim for computing server support, and J. Park, D. Bae, J. Lee, S. Kim, and J. Jung for their insightful discussions. This work was supported by the National Research Foundation of Korea (Grants No. RS-2023-NR077280, No. RS-2023-NR119928, and No. RS-2024-00413957).
Appendix A: Dependence of $D_{\rm eff}$ on $v_{0}$
@@
$g\rho(r)+V(r)+\tfrac{1}{2}mv(r)^{2}$ $=g\rho_{0}+\tfrac{1}{2}mv_{0}^{2},$ (12a)
$\nabla\!\cdot[\rho(r)v(r)]$ $=0.$ (12b)
@@
For a Gaussian obstacle $V(\mathbf{r})=V_{0}\exp(-2r^{2}/\sigma^{2})$, the solution for Eq. (14) is given by
@@
Appendix B: Dipole emitting frequency
@@
In Fig. 9(a), we show our numerical results of $f_{\mathrm{D}}$ for different obstacle strengths with $\sigma/\xi=20$. The drag frequency is well described by the linear relationship with $v_{0}$, but the critical velocity is slightly higher than that determined by the linear extrapolation of the emitting frequency. Noticeably, the value $a$ decreases with an increase of $V_{0}$, which is in stark contrast to the experimental observation. We attribute this discrepancy to the finite-length driving protocol used in the experiments: the limited obstacle motion introduces acceleration phases and interactions between emitted vortices, which can affect the measured shedding frequency. We confirmed this interpretation with our various numerical simulations.
The relationship of $f_{\mathrm{D}}=a(v_{0}-v_{c})$ suggests that the Strouhal number at $v_{0}=v_{\mathrm{th}}$ is given by
@@
where $E_{\mathrm{v}}$ is the energy cost of a vortex dipole. In Fig. 9(b), we show the variations of $E_{\mathrm{v}}=v_{0}F_{\mathrm{D}}/f_{\mathrm{D}}$ for different obtacle conditions. $E_{\mathrm{v}}$ increases steadily with the background flow speed $v_{0}$. Since the energy of a pair of vortices of opposite-signs separated by $D_{\mathrm{pair}}$ is given by $E_{\mathrm{v}}\propto\ln\!\left(D_{\mathrm{pair}}/\xi\right)$, the increase in $E_{v}$ implies that the separation $D_{\mathrm{pair}}$ increases with the velocity. This behavior is fully consistent with the $v_{0}$-dependence of the effective obstacle diameter $D_{\mathrm{eff}}$ (Appendix A).
@@
Appendix C: Extension to different obstacles
@@
For the impenetrable Gaussian obstacle (Figs. 10(a) and 10(c)), the shedding behavior and wake transition reproduce Ref. [25]: a von Kármán street of same-sign vortex pairs (K2 regime) for $\mathrm{Re}_{\rm s}<0.7$, followed by irregular shedding beyond $\mathrm{Re}_{\rm s}\simeq 0.7$. For a consistent comparison with Ref. [26], we analyzed the drag coefficient by subtracting the large-$\mathrm{Re_{s}}$ asymptote, $C_{\mathrm{D0}}$, rather than fitting $C_{\text{D}}$ directly. This yields $C_{\mathrm{D0}}=2.06$, with power-law exponents $\alpha_{1}=1.37$ for $\mathrm{Re_{s}}<0.7$ and $\alpha_{2}=1.55$ for $\mathrm{Re_{s}}>0.7$. Compared with the results of $C_{\mathrm{D0}}=2.37$, $\alpha_{1}=1.06$ and $\alpha_{2}=0.72$ in Ref. [26], our exponents are larger, likely because we use the flow-defined $D_{\mathrm{eff}}$, which increases with an increase of $v_{0}$, rather than a fixed Thomas-Fermi zero-density size as the characteristic length.
For the penetrable super-Gaussian obstacle (Figs. 10(b) and 10(d)), whose top is flatter than a standard Gaussian, the shedding dynamics closely mirror those of the penetrable Gaussian: periodic dipole emission at lower $\mathrm{Re_{s}}$ values, evolving into irregular shedding with same-sign vortex clusters at higher $\mathrm{Re_{s}}$ values. These results support $D_{\rm eff}$ as a robust, obstacle-agnostic scale that organizes superfluid wake dynamics.

### v2

@@ IV CONCLUSION AND OUTLOOK
We have shown that vortex shedding past a penetrable obstacle in a two-dimensional superfluid can be organized by a flow-defined length scale, the effective diameter $D_{\mathrm{eff}}$. Defined as the transverse extent of the region where the compressible component of the superflow exceeds the local speed of sound, $D_{\mathrm{eff}}$ delineates the Mach-1 contour that marks the onset of compressibility-driven instability. $D_{\mathrm{eff}}$ expands with $v_{0}$ and quantitatively matches the observed width of the vorticity distribution, confirming that the dynamically relevant scale is that felt by the flow rather than the geometric obstacle size. When forces and frequencies are normalized by $D_{\mathrm{eff}}$, the Strouhal and drag coefficients collapse onto universal curves versus the superfluid Reynolds number $\mathrm{Re}_{s}$, with a clear transition at $\mathrm{Re}_{s}\approx 2$. Thus, $D_{\mathrm{eff}}$ provides a unified similarity length governing both the shedding dynamics and mean dissipation in penetrable obstacle flows.
To test the broader applicability of our approach, we have also examined representative cases of impenetrable obstacles and non-Gaussian penetrable obstacles (Appendix D) and found scaling behavior similar to that for penetrable Gaussian obstacles. These results suggest that, provided the obstacle generates a bounded supersonic region and sustains vortex shedding, the effective-diameter framework remains useful over the parameter ranges examined here. The observed collapses further indicate that the Mach-1 contour provides a practical and physically motivated definition of the relevant obstacle length scale.
Nevertheless, it is important to note that the framework has an intrinsic limitation: it breaks down for $v_{0}/c_{\mathrm{s0}}>1$, where the flow becomes globally supersonic and no bounded Mach-1 region can be identified. In that regime, dissipation is expected to be governed by shock dynamics rather than vortex shedding [29, 53, 54], and a shock-based length scale may be relevant. In classical compressible flows, a shock layer controls the pressure drag past a bluff body [55].
The analogy between penetrable obstacles in a superfluid and classical porous bodies is worth noting. Porous obstacles permit partial through-flow, and their wake dynamics have been studied extensively [56, 57, 58]. In classical flows, the through-flow weakens the recirculation zone where vorticity accumulates and pushes it downstream, eventually suppressing vortex shedding. In our superfluid system, the locally supersonic region and the near wake play a qualitatively analogous role, and increasing penetrability similarly diminishes vortex shedding by allowing greater flow penetration. However, the underlying mechanisms are distinct: vorticity in porous-body wakes is generated in viscous shear layers, whereas in superfluids it arises from the breakdown of superfluidity within the locally supersonic region.
@@
Finally, we note that the universality revealed here can be tested experimentally. The shedding frequency or $\mathrm{St}$ can be measured directly using a conventional time-of-flight imaging, as shown in Refs. [19, 26, 28]. In addition, the drag force can be measured without direct mechanical probes using a control-volume analysis, as suggested in Ref. [25]. Moreover, a recent proposal based on matter–wave optics [59] suggests that both the superfluid velocity fields and the phase of the order parameter can be measured using phase–imprinting techniques [60, 61, 62], making it feasible to visualize the Mach-1 contour and measure $D_{\mathrm{eff}}$ directly. Such experimental methods would allow testing the predicted universal behaviors.
V acknowledgements
@@
We thank Haneul Kwak for sharing the simulation code, Myeonghyeon Kim for computing support, and J. Park, D. Bae, J. Lee, S. Kim, and J. Jung for insightful discussions. This work was supported by the National Research Foundation of Korea (Grants No. RS-2023-NR077280, No. RS-2023-NR119928, and No. RS-2024-00413957).
VI Data availavility
The data that support the findings of this article are openly available [63].
Appendix A: Details of dipole–to–cluster transition
Figure 8: Sequential snapshots of vortex nucleation near the obstacle for cases below and above the threshold $\mathrm{Re}_{\mathrm{s}}\approx 2$. (a–e) $\mathrm{Re}_{\mathrm{s}}=1.8$ $(v_{0}/c_{\mathrm{s}0}=0.31)$ and (f–j) $\mathrm{Re}_{\mathrm{s}}=4.4$ $(v_{0}/c_{\mathrm{s}0}=0.4)$, with $V_{0}/\mu=0.9$ and $\sigma/\xi=20$. Each column shows successive stages of wake development: before vortex nucleation [(a),(f)], nucleation event [(b),(g)], separation of opposite-sign vortices [(c),(h)], near-wake evolution [(d),(i)], and the start of the next cycle [(e),(j)]. The gray dashed line marks the Mach-1 contour.
In Sec. III A, we observed a sharp transition in the lift force near $\mathrm{Re}_{\mathrm{s}}\approx 2$, coincident with the emergence of same-sign vortex clusters. Figure 8 shows that vortices are nucleated by symmetric dipole pair production both below $\mathrm{Re}_{\mathrm{s}}<2$ [Figs. 8(a)8(e)] and above $\mathrm{Re}_{\mathrm{s}}>2$ [Figs. 8(f)8(j)]. The transition therefore reflects a change in near-wake dynamics, not in nucleartion mechanisms.
As $\mathrm{Re}_{\mathrm{s}}$ increases, the pair-production rate and near-wake vortex density increase, strengthening vortex–vortex interactions and thus amplifying small asymmetries. This favors an alternating wake mode rather than symmetric dipolar emission, qualitatively analogous to the onset of the von Kármán instability in classical bluff-body flow, where a symmetric wake gives way to an alternating vortex street above a critical Reynolds number [1]. Consistently, symmetry breaking is already visible during pair production in Fig. 8(j), where the positive vortex separates more clearly than the negative one. These observations suggest that the transition near $\mathrm{Re}_{\mathrm{s}}\approx 2$ is the onset of a wake instability that drives the downstream reorganization of emitted vortex dipoles into same-sign clusters.
By contrast, for an impenetrable obstacle, vortices of opposite signs are nucleated on opposite sides of the obstacle, so the wake naturally develops an alternating, or V-shaped, arrangement of vortex dipoles even at low velocities [17, 21]. With the symmetry broken from the outset, clustering manifests as a smoother crossover in the force response [39].
Appendix B: Dependence of $D_{\rm eff}$ on $v_{0}$
@@ VI Data availavility
$g\rho(\mathbf{r})+V(\mathbf{r})+\tfrac{1}{2}mv(\mathbf{r})^{2}$ $=g\rho_{0}+\tfrac{1}{2}mv_{0}^{2},$ (12a)
$\nabla\!\cdot[\rho(\mathbf{r})v(\mathbf{r})]$ $=0.$ (12b)
@@
For a Gaussian obstacle $V(\mathbf{r})=V_{0}\exp(-2r^{2}/\sigma^{2})$, we use the polar coordinate $(r,\theta)$ defined as $x=r\cos{\theta}$ and $y=\sin{\theta}$. The solution for Eq. (14) is given by
@@
Appendix C: Dipole emission frequency
@@
In Fig. 10(a), we show our numerical results of $f_{\mathrm{D}}$ for different obstacle strengths with $\sigma/\xi=20$. The drag frequency is well described by the linear dependence on $v_{0}$, but the critical velocity is slightly higher than that determined by the linear extrapolation of the emitting frequency. Noticeably, the value $a$ decreases with an increase of $V_{0}$, which is in stark contrast to the experimental observation. We attribute this discrepancy to the finite-length driving protocol used in the experiments: the limited obstacle motion introduces acceleration phases and interactions between emitted vortices, which can affect the measured shedding frequency. Additional numerical simulations support this interpretation.
The relation $f_{\mathrm{D}}=a(v_{0}-v_{c})$ suggests that the Strouhal number at the threshold velocity, $v_{0}=v_{\mathrm{th}}$, is given by
@@
where $E_{\mathrm{v}}$ is the energy cost of generating a vortex dipole. In Fig. 10(b), we show the variations of $E_{\mathrm{v}}=v_{0}F_{\mathrm{D}}/f_{\mathrm{D}}$ for different obstacle conditions. $E_{\mathrm{v}}$ increases steadily with the background flow speed $v_{0}$. Since the energy of a pair of vortices of opposite-signs separated by $D_{\mathrm{pair}}$ is given by $E_{\mathrm{v}}\propto\ln\!\left(D_{\mathrm{pair}}/\xi\right)$, the increase in $E_{v}$ implies that the separation $D_{\mathrm{pair}}$ increases with the velocity. This behavior is fully consistent with the $v_{0}$-dependence of the effective obstacle diameter $D_{\mathrm{eff}}$ (Appendix B).
@@
Appendix D: Extension to different obstacles
@@
For the impenetrable Gaussian obstacle (Figs. 11(a) and 11(c)), the shedding behavior and wake transition reproduce Ref. [24]: a von Kármán street of same-sign vortex pairs (K2 regime) for $\mathrm{Re}_{\rm s}<0.7$, followed by irregular shedding beyond $\mathrm{Re}_{\rm s}\simeq 0.7$. For a consistent comparison with Ref. [25], we analyzed the drag coefficient by subtracting the large-$\mathrm{Re_{s}}$ asymptote, $C_{\mathrm{D0}}$, rather than fitting $C_{\text{D}}$ directly. This yields $C_{\mathrm{D0}}=2.06$, with power-law exponents $\alpha_{1}=1.37$ for $\mathrm{Re_{s}}<0.7$ and $\alpha_{2}=1.55$ for $\mathrm{Re_{s}}>0.7$. Compared with the results of $C_{\mathrm{D0}}=2.37$, $\alpha_{1}=1.06$ and $\alpha_{2}=0.72$ in Ref. [25], our exponents are larger, likely because we use the flow-defined $D_{\mathrm{eff}}$, which increases with an increase of $v_{0}$, rather than a fixed Thomas-Fermi zero-density size as the characteristic length.
For the penetrable super-Gaussian obstacle (Figs. 11(b) and 11(d)), whose top is flatter than a standard Gaussian, the shedding dynamics closely mirror those of the penetrable Gaussian: periodic dipole emission at lower $\mathrm{Re_{s}}$ values, evolving into irregular shedding with same-sign vortex clusters at higher $\mathrm{Re_{s}}$ values. These results support $D_{\rm eff}$ as an obstacle-independent scale that organizes superfluid wake dynamics.